OpenStudy (anonymous):

Integral from 0 to 2 of: (x^3+2x^2-5x-1) / (sqrt x)

OpenStudy (anonymous):

where are you stuck?

OpenStudy (anonymous):

I know to do this: \[\frac{ x^4 }{ 4 } + \frac{ 2 }{ 3 }x^3 -\frac{ 5 }{ 2 }x^2 -x\] to the top

OpenStudy (anonymous):

And i know the bottom is just \[x ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

NOOO!!! can't be it

OpenStudy (anonymous):

do you multiply the top by x^2 first?

OpenStudy (anonymous):

The integrand only: \[\dfrac{x^3+2x^2-5x-1}{x^(1/2)}=(x^3+2x^2-5x-1)(x^{-1/2})\] distribute it, then put the integral sign in, take it as usual.

OpenStudy (anonymous):

can you step up?

OpenStudy (anonymous):

I know how to distribute that, but how do you know to switch the exponent to a negative 1/2?

OpenStudy (anonymous):